In the 2005 Major League Baseball season, the Chicago White Sox won 27 fewer than twice as many games as they lost. They played 162 regular season games. How many wins and losses did the team have?
Found 2 solutions by checkley71, ptaylor:
Answer by checkley71(8403)

WINS=2L-27
(2L-27)+L=162
3L=162+27
3L=189
L=189/3
L=63 NUMBER OF LOSSES.
W=2*63-27=126-27=99 NUMBERS OF WINS.
PROOF
63+99=162
162=162

Answer by ptaylor(2198)

Lex x=number of games they lost
Then 2x-27=number of games they won
Number of games they lost plus number of games they won = 162, so:
x+2x-27=162 add 27 to both sides
x+2x-27+27=162+27 combine like terms
3x=189 divide both sides by 3
x=63---------------------------number of games they lost
2x-27=2*63-27=126-27=99 ------------------number of games they won
CK
63+99=162
162=162
Also 99 is 27 less than two times 63
Hope this helps---ptaylor